OpenStudy (anonymous):

x^3-1=0 I understand how one of the answers is 1, but why is it also -1+-root 3 all over 2

OpenStudy (anonymous):

it's not. 1 is the only root

OpenStudy (anonymous):

no i don't think so...-1+-root 3 all over 2 is listed in the answer key

OpenStudy (anonymous):

then either there is typo or you're looking at the solution to a different problem. 1 is the only solution.

OpenStudy (anonymous):

Look at the solution again, is there an "i" , the imaginary number, anywhere?

OpenStudy (anonymous):

no, but do you think there should be? its possible he made a typo

OpenStudy (anonymous):

1 is the real root. the other two complex root are \[\frac{-1 \pm \sqrt{3}\ i}{2}\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

(x^3 - 1) = (x - 1) (x^2 + x + 1) the complex roots are found by solving x^2 + x + 1 using quadratic fomula

OpenStudy (anonymous):

thanks!